Home Sweet Home, a 50-room hotel, currently operates at 52.00% occupancy. Total rooms revenue generated from the room sales is $15,600 with a marginal cost of $90.00 per room. The management is considering discounting the rack rate by 15.00% so that they can increase the occupancy. What equivalent occupancy percentage must be achieved to maintain the same levels of total room contribution margin level per room with the discount in the rack rate (assume that marginal cost will remain the same for the discount situation)? Group of answer choices 63.14% 75.56% 68.70% 71.30%
Answer #1

Current Scenario

Revenue from room sales = $ 15,600

Occupancy rate = 52%

Number of rooms occupied = 50*52% = 26 rooms

Room rack rate = 15,600/26 = $ 600 per room

Room rack rate

$ 600

Marginal cost

($ 90)

Contribution per room

$ 510

Contribution margin ratio = 510/600 = 85%

Under new scenario, the discount rate of 15% is considered.

Room rack rate

$ 600

Less: Discount at 15%

($ 90)

$ 510

Less : Marginal cost

($ 90)

Contribution per room

$ 420

Contribution margin = 420/600 *100 = 70%

Therefore,

The equivalent occupancy would be

                             = Current Occupancy % * Current contribution margin/New Contribution margin

                                = 52% * 85%/70%

                                = 63.14%

Hence, the required occupancy percentage to maintain the same levels of total contribution margin level would be 63.14%

Kindly upvote